OpenStudy (sh3lsh):

Multivariable: Assume there is an opaque ball of radius 1 centered at the origin. Suppose that you stand at the point (2,3,0) and look in the direction of a point that is not visible because it is behind the ball. You will then be looking at a point on the sphere. Find the point on the sphere at which you look if you are looking in the direction of (−2,−3,2).

OpenStudy (thomas5267):

I would first find the line connecting the point you stand and the point you look. Then find the intersection between the point and the sphere.

OpenStudy (thomas5267):

*the line

OpenStudy (sh3lsh):

So I found the vector from PR, P being (2,3,0) and R from (-2,-3,2). It's just <-4,-6,2> but I'm unsure how to Incorporate that now.

OpenStudy (thomas5267):

You have found the direction vector of the line. Try find the equation of the line.

OpenStudy (sh3lsh):

I have a line now of (2,3,0) + (-4,-6,2)t. The equation of the sphere is just (x-1)^2 + (y-1)^2 + (z-1)^2 = 1, correct? Should I set the x,y,z equations equal to each other to get a parameter t, and plug it in?

OpenStudy (sh3lsh):

Hmmm, how would I go about setting the equations equal to each other? Parameter the sphere? If so, how would I go about that?

OpenStudy (thomas5267):

I think the equation of sphere that is centred at origin is \(x^2+y^2+z^2=1\). Then you would write the linear equation as \(x=\text{sth},\,y=\text{sth},\,z=\text{sth}\) and substitute back into the sphere equation.

OpenStudy (sh3lsh):

So, I just substituted the corresponding directions into the sphere equation. (2-4t)^2+(3-6t)^2+(2t)^2=1 I just put this into Wolfram to get the solutions and I got, 3/7 and 1/2. I used the smaller one, because that's the one I'll see. I plugged this back in, and the solutions work! Thanks so much!

OpenStudy (thomas5267):

Wolfram is nice, but make sure you can solve it by hand or you will be screwed in an exam.

OpenStudy (sh3lsh):

Hahaha, that's very true. I hadn't thought of that!

OpenStudy (thomas5267):

I got over-reliant on Wolfram Alpha and screwed up a mock exam. Glad that that was a mock.